OpenStudy (cookie_2046):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 12, c = 19 A. 14.03 B. 14.38 C. 14.73 D. 15.07

OpenStudy (anonymous):



OpenStudy (anonymous):

familiar with Pythagorean?

OpenStudy (cookie_2046):

Wait

OpenStudy (cookie_2046):

is the answer going to be c?

OpenStudy (cookie_2046):

Use the Pythagorean Theorem. a^2 + b^2 = c^2 a^2 + 12^2 = 19^2 Simplify the exponents: a^2 + 144 = 361 Subtract 144 to both sides: a^2 = 217 Find the square root of both sides: a = 14.73

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

good job

OpenStudy (cookie_2046):

Yay so then i'm correct :D

OpenStudy (anonymous):

yep